OpenStudy (anonymous):

suppose that f(x) =k(6)^X. If f(-2), then f(3) = ?

OpenStudy (anonymous):

if f(-2)= what?

OpenStudy (anonymous):

oh sorry! f(-2) = 5

OpenStudy (anonymous):

is this \[f(x)=k\times 6^x\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then we can solve for k by knowing \[5=k\times 6^{-2}=\frac{k}{36}\] making \[k=180\] i guess

OpenStudy (anonymous):

so you have \[f(x)=180\times 6^x\] and so \[f(3)=180\times 6^3\] whatever that is

OpenStudy (anonymous):

38880